OpenStudy (mb194227):

Please help me solve x^3+12x^2+47x-60=0.

OpenStudy (mathmale):

Note that this is a 3rd order polynomial, so the quadratic equation does not apply here. I would suggest you either try to factor, or try to identify possible roots by looking at all factors of -60.

OpenStudy (mathmale):

Factors of -60 include plus or minus 1, p or m 2, p or m 3, p or m 4, and so on. Do you have any experience with synthetic division? If so, use it here to determine whether or not 4 is a root o the given cubic equation.

OpenStudy (eliesaab):

Try to see that x=1 is a root.

OpenStudy (eliesaab):

It is clear that 1+12+47-60=60-60=0

OpenStudy (eliesaab):

Divide your polynomial by x-1

OpenStudy (eliesaab):

You get \[ \frac{x^3+12 x^2+47 x-60}{x-1}=x^2+13 x+60 \]

OpenStudy (eliesaab):

Use the quadratic formula to finish